OpenStudy (anonymous):

A dartboard consists of two concentric circles. The probability of hitting the inner circle is 16%. If the outer circle has a diameter of 20 inches, what is the radius of the inner circle?

OpenStudy (anonymous):

5:) because the radius to the big circle is the diameter to the samll one. Take half of the small diameter and u get 5... hope this helps:)

OpenStudy (anonymous):

The answer choices are 2,4,8 and 16

OpenStudy (anonymous):

ok wait i totally messed tht up...SORRY

OpenStudy (anonymous):

ok sry maybe u should take @jim_thompson5910 help... sorry

jimthompson5910 (jim_thompson5910):

You first need to find the area of the larger circle. The larger circle has a diameter of 20, or a radius of 20/2 = 10 So what is the area of this larger circle?

OpenStudy (anonymous):

100pi?

jimthompson5910 (jim_thompson5910):

good, so if the probability of landing in the smaller circle is 16%, then this means P(Landing in Smaller Circle) = (Area of Smaller Circle)/(Area of Larger Circle) 16/100 = x/(100pi) cross multiply and solve for x to get 16/100 = x/(100pi) 16*100pi = 100x 1600pi = 100x 100x = 1600pi x = 1600pi/100 x = 16pi Which means that the area of the smaller circle is 16pi square inches

jimthompson5910 (jim_thompson5910):

now that you know the area of the smaller circle is 16pi, you can use this to find the radius of the inner circle

jimthompson5910 (jim_thompson5910):

how do we do this?

OpenStudy (anonymous):

divide it by 2?

jimthompson5910 (jim_thompson5910):

Remember that A = pi*r^2 is the area of a circle

jimthompson5910 (jim_thompson5910):

So A = pi*r^2 16pi = pi*r^2 16 = r^2 ... Divide both sides by pi. r^2 = 16 r = ???

OpenStudy (anonymous):

r=4?

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

Thank you :D

jimthompson5910 (jim_thompson5910):

np